OpenStudy (anonymous):

How to solve 1/1+1/2=1/x?

OpenStudy (anonymous):

I realize this is quite basic but I'm stuck on the first step. I had to make this equation based on a word problem, and I got this far, so what else do I do?

OpenStudy (potatoes.ramu):

x =2/3

OpenStudy (potatoes.ramu):

1+1/2= 3/2

OpenStudy (anonymous):

I know what x equals, but how do I get there? Potatoes.ramu was right.

OpenStudy (potatoes.ramu):

3/3=1/x

OpenStudy (potatoes.ramu):

3/2=1/x sorry typo

OpenStudy (anonymous):

How did you get 3/2?

OpenStudy (potatoes.ramu):

Once you cross multiply, x= 2/3

OpenStudy (potatoes.ramu):

Add 1+1/2.. you get 3/2 :)

OpenStudy (anonymous):

Oh, I get it now! Thank you @Potatoes.ramu!!!

OpenStudy (potatoes.ramu):

No problem! :) Best of luck! :)